### v1

Abstract
Nonlinearity is often associated with amplification, instability, or emergent behavior — where small inputs produce disproportionately large outputs. Yet in many physical systems, nonlinearities act as limiting mechanisms, suppressing the response below linear expectations, as exemplified by the breakdown of Hooke’s law at large forces. These limiting effects can become significant in the optical control of quantum materials, where large-amplitude lattice motion is essential for reaching functional states far from equilibrium. Displacive excitation of coherent phonons, perhaps the most widely used control strategy, typically assumes linearity between the driving force on the lattice and the optically excited carrier density. However, the breakdown of this assumption remains largely unexplored, and strategies are needed to overcome the resulting nonlinear bottlenecks. Here, we show that sequential optical excitation overcomes the nonlinear amplitude saturation of a key phonon mode in the van der Waals ferroelectric Td-WTe2. Using time-resolved second harmonic generation, we track the response of the interlayer sliding mode governing ferroelectricity to both single- and double-pulse excitation. With increasing optical excitation, the phonon amplitude saturates and declines — a limitation we trace to band-specific electron–phonon coupling that weakens the driving force on the nuclei. Sequential in-phase excitation of coherent sliding motion, which outlasts the electronic excitation, avoids population of counteracting electronic states. This enables higher phonon amplitudes for the same optical energy input. We exploit the enhanced lattice response to perform high-amplitude vibrational spectroscopy in the excited state, revealing a novel form of anharmonic phonon coupling that emerges exclusively under nonequilibrium conditions, and previously elusive coherent vibrational modulations of electronic states. Our results establish a more general picture of electron-phonon and phonon-phonon coupling far from equilibrium, opening new possibilities for ultrafast ferroic switching and lattice control in quantum materials.
keywords:
Ultrafast dynamics, nonlinear phononics, nonlinear optics, ferroelectrics
Large-amplitude coherent vibrational motion is a powerful route to accessing nonequilibrium states of matter[1, 2, 3, 4, 5, 6, 7]. In nonlinear phononics [3], intense mid-infrared fields resonantly drive infrared-active phonons to explore the potential energy surface and trigger transitions into metastable ferroelectric [8], magnetic [9], or electronic phases [10, 11]. This approach often relies on anharmonic coupling between the driven mode and Raman-active lattice distortions associated with phase transitions. Alternatively, direct excitation of Raman-active modes via displacive excitation of coherent phonons (DECP) [12, 13] exploits the coupling between transiently populated electronic states and targeted lattice modes. At the same time, decisively large vibrational amplitudes often remain elusive in DECP due to damage thresholds or anharmonic phonon decay in highly excited states [14]. For vibrational coherence outlasting electronic excitation, sequential excitation offers a promising strategy to overcome these limitations [15, 16, 17]. However, how far sequential excitation can extend the boundaries of vibrational control under intrinsic constraints remains an open question.
Here, we demonstrate that splitting the energy of an optical pulse into two temporally separated excitations overcomes amplitude saturation of the polar phonon mode in a van der Waals ferroelectric by avoiding the population of counteracting electronic states, thereby enabling access to nonlinear lattice phenomena. Using time-resolved second harmonic generation (trSHG) [18, 19] and ab initio calculations, we investigate the coupling between interlayer sliding motion and ferroelectric polarization, and trace the amplitude limit of the coherent sliding phonon to the occupation of high-lying electronic states. Re-excitation of vibrational coherences following relaxation of the electronic subsystem circumvents this saturation threshold and amplifies sliding motion beyond the single-pulse limit. Tracking the enhanced lattice response reveals a strong correlation between sliding mode amplitude, damping, and frequency, which serve as direct signatures of a new form of anharmonic coupling directly mediated by the coherent nuclear motion. At highest sliding amplitudes, a modulation of the trSHG signal at twice the sliding mode frequency signals a strong, lattice-driven modulation of electronic states and nonlinear optical properties—an effect previously obscured by limited vibrational amplitudes.
As a model system, we investigate WTe2, a ferroelectric van der Waals semimetal that has recently attracted attention for its giant, non-saturating magnetoresistance and ultrafast optical switching between topologically distinct states [20, 21, 22, 23, 24, 25, 26]. In its thermodynamically stable ground state, the Td polymorph (stable for temperatures $T<565\,\text{K}$), WTe2 adopts a trilayer structure that breaks inversion symmetry perpendicular to the van der Waals planes (see Fig. 1b). An uncompensated interlayer vertical charge transfer gives rise to a ferroelectric polarization oriented along the crystallographic $c$-axis, that is, normal to the layers [27, 28, 29, 30]. Ultrafast optical excitation across a broad spectral range (from terahertz to visible) depopulates antibonding orbitals between neighboring layers [21, 25, 24, 31]. This launches an interlayer sliding mode at $f_{\text{s}}=0.24\,\text{THz}$, which pushes the system toward a high-symmetry configuration on the potential energy surface restoring inversion symmetry [21, 25, 32, 33, 34, 35] (Figs. 1b and c). Interlayer sliding displacements are expected to strongly affect the material’s ferroelectric properties on ultrafast time scales [29], though the precise interplay between structural dynamics and polarization deserves further consideration.
We use symmetry considerations to investigate the connection between structural distortions and ferroelectric polarization in WTe2. The polar Td phase belongs to the non-centrosymmetric space group $Pmn2_{1}$, while the non-polar 1T’ phase crystallizes in $P2_{1}/m$. Since these two space groups are not in a group–subgroup relation, a direct transition between them cannot be achieved by a single phonon mode. Instead, we identify the true high-symmetry reference structure in the centrosymmetric $Pnma$ space group (see Fig. 1b). From this phase, two distinct lattice instabilities emerge. First, the interlayer sliding mode ($A_{1}$$\Gamma_{2}^{-}$ irreducible representation, $f_{\text{s}}=0.24$ THz), which breaks inversion symmetry, lowers the structure to $Pmn2_{1}$, and directly controls the ferroelectric polarization $P_{z}$. Second, an intralayer shear mode ($B_{1}$$\Gamma_{3}^{+}$, $f=3.4$ THz), which transforms like the $\varepsilon_{yz}$ component of the elastic tensor, lowers the symmetry to $P2_{1}/m$, and governs monoclinicity. Consequently, the Td and 1T’ structures are connected only via a two-step pathway: excitation of the interlayer sliding mode to reach the $Pnma$ reference structure, followed by intralayer shear distortion leading to 1T’. In the Td phase, both modes influence polarization dynamics through
$F=F_{Td}+\omega_{A1}^{2}(P-P_{z})^{2}+\omega_{B1}^{2}Q_{B1}^{2}+\gamma(P-P_{z})Q_{B1}^{2},$ (1)
where $F$ is the free energy and $P_{z}\sim Q_{A1}$. In other words, the interlayer sliding couples directly to $P_{z}$, while intralayer shearing couples indirectly through nonlinear phonon-phonon interactions.
Second harmonic generation, owing to its inherent sensitivity to symmetry breaking, is an ideal probe of structural degrees of freedom governing ferroelectric order [36]. Tracking the transient symmetry changes in Td-WTe2 following single-pulse optical excitation via polarization-resolved trSHG (Fig. 1a) reveals a rapid initial drop in SHG intensity within 500 fs, concomitant with prominent oscillations at the interlayer sliding mode frequency (see Figs. 1d). While all accessible tensor elements oscillate coherently at the same frequency, a strong initial suppression is observed exclusively in the vertical lobe associated with the $\chi_{zyy}$ component (compare frames at $0.5\,\text{ps}$ and $0.5\,\text{ps}$ in Fig. 1d; for further details, see Fig. S2). We speculate that this anisotropic response results from an additional structural distortion along the $c$-axis, as suggested by density functional theory calculations under optical excitation [29].
Sliding mode amplitude limitation
To explore the limits of displacive sliding mode excitation, we perform fluence-dependent trSHG measurements (see Fig. 1e). Increasing the absorbed fluence from $0.4-6.5\,\text{mJ}\,\text{cm}^{-2}$ progressively shifts the equilibrium position $Q_{0}$ toward the centrosymmetric 1T’ configuration (Fig. 1c) and leads to a pronounced softening of the sliding mode, reflected in both long-lived signal suppression and fluence-dependent frequency shifts (Fig. 1g). Most notably, however, the sliding mode amplitude initially increases with fluence, then saturates around an absorbed fluence of $F_{\text{sat}}\approx 1.6\,\text{mJ}\,\text{cm}^{-2}$, and subsequently declines at higher excitation levels (see Fig. 1f). This behavior is consistent with observations from transient reflectivity and time-resolved X-ray diffraction measurements on Td-WTe2 [35] and the closely related compound Td-MoTe2 [37]. This amplitude saturation places a fundamental limit on the extent to which coherent sliding mode excitation can be used to control ferroelectricity.
Understanding the fluence-dependent behavior of the sliding mode amplitude requires going beyond the standard DECP model, which assumes a linear relationship between excited carrier density and coherent phonon amplitude [12]. Instead, we account for the contributions of individual electronic states at different energies to the effective electron–phonon coupling at the sliding mode frequency. To this end, we perform ab initio calculations [38, 39] to determine the driving force on the sliding mode as a function of the optically-induced transient electronic temperature (see Methods Section for details).
Figure 2a displays the state-resolved driving forces for the sliding mode at electronic temperatures $T_{\text{elec}}=2000\,\text{K}$ and $T_{\text{elec}}=4000\,\text{K}$, that is, above and below the saturation threshold. The contributions of individual states are weighted by their occupation, governed by changes $\Delta f_{\text{elec}}$ in the Fermi–Dirac distribution at the transient electronic temperature following optical excitation. At low excitation laser fluences, that is, low electronic temperatures, changes in occupation of electron- and hole-like states near the Fermi energy $E_{\text{F}}$ generate net positive sliding forces. With increasing fluence, higher-energy states become populated, which instead counteract sliding mode excitation.
In a real-space picture [40], the generation of photo-excited electrons and holes near $E_{F}$ induces a spatial modulation of the charge density within the unit cell. Through Coulomb interactions, this transient distribution drives nuclear displacements aligned with the sliding mode pattern. In contrast, population of high-energy states produces a more homogeneous charge density, reducing net Coulomb forces on the nuclei and thereby suppressing sliding-mode-driven lattice motion (see fluence-dependent maximum sliding amplitude $Q_{\text{max}}(F)$ in Fig. 2b (violet circles)). The saturation behavior is unique to the sliding mode, whereas higher-frequency modes exhibit a continuous, monotonic increase in amplitude. Comparison between experimental data (violet triangles) and the ab initio model (violet circles) reveals close agreement in the fluence-dependence of the sliding mode amplitude (Fig. 2c). More specifically, our calculations predict a saturation of the sliding mode amplitude at an electronic temperature of $T_{\text{sat}}=3100\,\text{K}$. In experiment, this corresponds to an absorbed fluence of $F_{\text{sat}}=1.6\,\text{mJ}\,\text{cm}^{-2}$, consistent with the experimental value. Beyond this fluence, both experiment and theory show a decline in sliding mode amplitude.
Coherent control and amplification of sliding motion
The amplitude limitation restricts coherent vibrational manipulation of the material’s ferroelectric [27, 20, 29] and topological [21, 24] properties to low fluences. Yet it is the high-fluence regime, where the system approaches the centrosymmetric configuration, that offers the opportunity to harness phonon coherences for coherent on–off switching of ferroelectricity and manipulation of Weyl nodes. This motivates the search for control schemes that minimize electronic excitation while maximizing achievable phonon amplitudes. Repeated in-phase excitation of coherent phonons by optical pulse trains has been shown to selectively amplify targeted vibrational modes to high amplitudes [15, 16]. A necessary condition for such a ‘cold’ driving scheme is vibrational coherence outlasting the electronic excitation—a criterion met in WTe2, as confirmed by our results and previous time-resolved studies [34, 35, 21, 25, 32, 33].
We therefore aim to control the sliding mode amplitude by a double-pulse excitation scheme (Fig. 3a) [16, 17, 41, 42, 43]. To this end, we split our pump pulses in two and record trSHG traces as a function of the mutual delay $\Delta t_{\text{p-p}}$ between the two pump pulses and the double-pump-probe ($\Delta t_{\text{p-pr}}$) delay (see Fig. 3b). Strong modulations of the vibrational amplitude for different double-pulse delays $\Delta t_{\text{p-p}}$ demonstrate coherent control over the sliding mode amplitude $A_{\text{s}}$ in the excited state (see selected SHG traces in Fig. 3c), corroborated by ab initio calculations (Fig. 3d). Comparing double-pulse SHG traces for $\Delta t_{\text{p-p}}=0$, where both pump pulses act as a single excitation, and $\Delta t_{\text{p-p}}=2T_{\text{s}}$ ($T_{\text{s}}=2\pi/f_{\text{s}}$), we find that, for equal total optical energy, sequential optical excitation yields up to 50 % higher sliding amplitudes (see Fig. 3e). This demonstrates that sequential excitation enables larger vibrational amplitudes in the targeted nonequilibrium state than those attainable with single-pulse excitation at the same total fluence (see Fig. 3g).
Analyzing the sliding mode amplitude after the second excitation as a function of $\Delta t_{\text{p-p}}$ reveals a strong suppression for overlapping pump pulses compared to the delay-dependent amplitude expected in coherent control experiments (see red-shaded area near $\Delta t_{\text{p-p}}=0$ in Fig. 3f). The suppression persists for inter-pulse delays up to $\sim$2 ps (compare violet trace in Fig. 3f). This timescale agrees with the carrier lifetimes determined by time- and angle-resolved photoemission spectroscopy (trARPES) [26], supporting our hypothesis that excitation of charge carriers to high energies above the Fermi level counteracts sliding mode generation. While saturable absorption in WTe2 on comparable timescales has been reported [44], the magnitude of that effect appears insufficient to explain the pronounced sliding mode saturation and, in particular, the subsequent suppression observed here.
Large-amplitude coherent vibrational spectroscopy
In a next step, we exploit sequential excitation to enable large-amplitude coherent vibrational spectroscopy in the excited state. In particular, double-pulse excitation not only facilitates enhanced vibrational amplitudes but also allows precise control over the amplitude while keeping other key degrees of freedom, such as electronic excitation and temperature, constant [45]. This makes it possible to isolate correlations between the coherent sliding mode amplitude on the one hand, and its damping and frequency on the other hand (see Fig. 4a) [46].
Using this approach, we find a strong correlation between the sliding mode amplitude $A_{\text{s}}$ and its damping constant $\gamma_{\text{s}}=\tau_{\text{s}}^{-1}$ across the entire amplitude range (Fig. 4a), indicating that the damping is directly mediated by coherent nuclear motion. Crucially, this decay channel is absent in equilibrium crystals, where nuclear wave packets remain centered at their equilibrium positions regardless of temperature. The observed amplitude-dependent damping therefore reflects higher-order anharmonic phonon–phonon interactions [46, 47] that emerge exclusively under non-equilibrium conditions.
The most probable relaxation pathway for the sliding mode involves a Klemens-like decay of the sliding mode into two lower-energy acoustic phonons with opposite momentum [48, 49, 50] (see Fig. 4c). Due to the low energy of the sliding mode ($\hbar\omega_{\text{ph}}\approx 1\,\text{meV}$), only acoustic modes near the Brillouin zone center are energetically compatible with such a down-conversion process. Amplitude-dependent damping traces recorded at two different combined fluences support this scenario (Fig. 4b). While the two traces exhibit the same linear correlation between amplitude and damping, the slopes $(\Delta\tau/\Delta A)_{\text{s}}$ increase with increasing fluence. We attribute this to a fluence-dependent transient temperature rise $\Delta T$, which enhances the thermal population of phonons according to the Bose–Einstein distribution. As anharmonic decay is a stimulated process [51], increased occupation of the daughter phonon modes leads to an elevated decay rate of the sliding mode. For the two absorbed fluence values of $F=2.2\,\text{mJ}\,\text{cm}^{-2}$ and $F=4.1\,\text{mJ}\,\text{cm}^{-2}$, we estimate corresponding temperature increases of $\Delta T=70\,\text{K}$ or $130\,\text{K}$, respectively (see Methods Section for details).
Similarly, the sliding mode frequency exhibits a clear anti-correlation with the sliding amplitude (compare red and brown traces in Fig. 4a): as the amplitude increases, the frequency softens. Since both the lattice temperature and the level of electronic excitation are fixed in double-pulse experiments, a significant contribution from electronic softening, as reported in other semimetals [52], can be excluded. We therefore attribute the observed amplitude-dependent frequency shift to intrinsic lattice anharmonicity, explored by large-amplitude coherent nuclear motion.
We note that these insights into the processes underlying coherent phonon decay cannot be straightforwardly obtained from conventional (fluence-dependent) optical pump–probe experiments alone, but require precise control over the phonon amplitude in a targeted excited state, as enabled by double-pulse excitation (compare pump-probe and pump-pump-probe data in Fig. 4b)).
In a last step, we consider the impact of large-amplitude sliding motion on the electronic structure. For the highest sliding amplitudes in the excited state, Fourier transforms of the time-resolved SHG traces show an additional contribution at twice the sliding mode frequency ($2\omega_{\text{s}}$, see Fig. 5). This component is absent for single-pulse excitation, for double-pulse excitation near temporal overlap (Fig. 5a-c), and in transient reflectivity studies [33, 32, 35, 34]. This suggests that its generation requires vibrational amplitudes beyond the sliding mode saturation threshold, achievable only through sequential excitation. Phonon harmonics have been attributed to anharmonicities in the interatomic potential [53], or to periodic modulation of electronic states through electron-phonon coupling [37, 54, 55]. In the first case, frequency doubling originates directly from anharmonicities in the atomic motion $Q(t)$, which lead to a second harmonic component in the lattice displacement. In the second case, the coherent lattice motion periodically modulates the material’s second-order susceptibility, $\chi^{(2)}=\chi^{(2)}(t)$, thereby generating sidebands in the SHG spectrum through dynamic electron–phonon coupling. This mechanism is commonly associated with deformation potential effects [37] or phonon Floquet phenomena [56], where the electronic structure is periodically driven by vibrational motion.
Although second harmonic phonon generation in the intrinsically anharmonic potential of WTe2 may contribute, several experimental observations point to a modulation of $\chi^{(2)}$ as the most likely origin of the $2\omega_{\text{s}}$ feature. First, recent transient reflectivity measurements on the related compound Td-MoTe2 have reported similar harmonic and sum-frequency components for probe photon energies near electronic resonances [37]. Second, trARPES has uncovered periodic shifts of electronic band energies in WTe2 following sliding mode excitation [26], corroborated by DFT calculations [35]. Together, these results suggest a scenario in which amplified lattice motion periodically modulates the energy of electronic states, thereby driving the nonlinear probing process in and out of resonance. The resulting amplitude modulation of the optical SHG response leads to the emergence of frequency sidebands (see derivation in the Methods section).
While phonon-induced energy shifts near the Fermi level are well established, our results show that repeated optical excitation can modulate optical transitions even in band structure regions that are typically only weakly coupled to structural modes. This capability is particularly relevant for the generation of optical frequency combs via coherent phonons [54, 55]. Our work extends this approach to a broader range of photon energies, beyond linear transmission or reflectivity, and to the regime of optical second-harmonic generation, offering an alternative pathway to the modulation of optical nonlinearity that is complementary to recent advances in Floquet engineering [18].
Discussion and summary
In summary, we have shown that redistributing the energy of a single ultrafast excitation into two or more pulses can overcome the amplitude limitation of key vibrational modes. Extending this concept to multi-pulse sequences of sub-threshold excitations may enable ’cold’ driving of Raman-active modes deep into the nonlinear regime [1, 15, 16]. By analogy with nonlinear phononics [3, 8, 9], the resulting large-amplitude motion along Raman-active coordinates could access novel dynamical phases of matter and facilitate ultrafast, energy-efficient switching of order parameters. Our approach may be particularly relevant for the class of sliding ferroelectrics, where—due to broken inversion symmetry—decisive modes are typically both Raman and IR-active but carry only small dipole moments, limiting the efficiency of direct resonant excitation.
While a linear relationship between electronic temperature and vibrational amplitude is a central assumption of the DECP model [12], there is no fundamental reason for it to hold universally. The principles demonstrated here for WTe2 should therefore be broadly applicable to other materials and their structural modes. In this context, high-amplitude vibrational spectroscopy enabled by sequential excitation [57, 46] offers a powerful tool to probe phonon–electron and phonon–phonon interactions in nonequilibrium states beyond conventional pump–probe schemes. Understanding how decisive coherent phonons decay under such conditions is essential for advancing vibrational control strategies. For WTe2, our work identifies a nonequilibrium decay channel directly governed by coherent nuclear motion, highlighting amplitude-dependent damping as a key factor in determining phonon lifetimes as well as energy and heat transfer. We envision that the enhancement of vibrational coherences near topologically trivial and paraelectric structural configurations opens new avenues for coherent control of both topology and ferroelectricity in functional materials.
Methods
Sample preparation. WTe2 single crystals (commercially ordered from HQ Graphene) were exfoliated to remove the oxidized surface layers and immediately mounted in an optical microscopy cryostat. The sample space was constantly purged with nitrogen gas to minimize oxidization of the pristine surface layers. Peak SHG and oscillatory signals from the sample showed less than $2\,\%$ degradation over several weeks. All experiments were carried out at room temperature.
Experimental setup. In single-pulse trSHG measurements, structural dynamics are initiated by focusing pump pulses (center wavelength $\lambda_{\text{p}}=1.5\,\mu\text{m}$, photon energy $\hbar\omega_{\text{p}}=0.83\,\text{eV}$, pulse duration $\Delta\tau_{\text{p}}=100\,\text{fs}$, absorbed fluence $F_{\text{p}}=0.4–6.5\,\text{mJ}\,\text{cm}^{-2}$, polarization along the $b$-axis of the WTe2 crystal, repetition rate $f_{\text{rep}}=1\,\text{kHz}$) onto the sample surface under normal incidence using a plano-convex lens ($f_{\text{pump}}=100\,\text{mm}$). The resulting dynamics are probed by monitoring the second harmonic signal generated by probe pulses (center wavelength $\lambda_{\text{pr}}=0.8\,\mu\text{m}$, photon energy $\hbar\omega_{\text{pr}}=1.55\,\text{eV}$, pulse duration $\Delta\tau_{\text{p}}=120\,\text{fs}$, fluence $F_{\text{p}}=2.8\,\text{mJ}\,\text{cm}^{-2}$, repetition rate $f_{\text{rep}}=1\,\text{kHz}$) as a function of the time delay $\Delta t_{\text{p-pr}}$ (mechanical delay stage: Physik Instrumente M-414.DG), and for different combinations of incident and detected probe polarizations. Probe pulses are focused onto the sample in the $ac$-plane and at a $45^{\circ}$ angle to the surface normal using a plano-convex lens ($f_{\text{pump}}=75\,\text{mm}$). For the measurements shown in Figs. 1- 5, SHG signals are acquired in a cross-polarized configuration, with the incident (detected) probe polarization aligned parallel (perpendicular) to the $ac$-plane of the crystal.
Parasitic SHG signals originating from optical elements in the probe path are suppressed using a long pass filter (Schott RG630, $2\,\text{mm}$). After reflection from the sample, second harmonic and fundamental beams are separated using a short pass filter (Schott BG39, $2\,\text{mm}$) and a bandpass filter (center wavelength $400\,\text{nm}$, bandwidth $25\,\text{nm}$). The SHG signal is collected by a plano-convex lens (focal length $f=75\,\text{mm}$), detected using a photomultiplier tube, and processed with electronic amplification and time-domain filtering via a boxcar integrator to enhance the signal-to-noise ratio. Each data point is averaged over 5000 pulses ($t_{\text{int}}=5$ s). The temporal overlap between pump and probe pulses (time zero) is determined by sum-frequency generation in a beta-barium-borate (BBO) crystal placed at the sample position. To ensure homogeneous excitation of the probed region, the pump spot size on the sample ($280\times 280\,\mu\text{m}$) is chosen significantly larger than the probe spot ($80\times 40\,\mu\text{m}$).
For double-pulse trSHG measurements, pump-pulse pairs are generated using a Mach–Zehnder-type interferometer and focused colinearly onto the sample. The relative delay $\Delta t_{\text{p1-p2}}$ between pulses p1 and p2 is controlled by a motorized delay stage (Physik Instrumente M-111.DG) in one interferometer arm. The pulse energies and polarizations of p1 and p2 are adjusted independently using neutral density filters and half-wave plates placed in each arm. Temporal overlap ($\Delta t_{\text{p1-p2}}=0$) is determined by measuring the SHG signal from the sample as a function of $\Delta t_{\text{p1-p2}}$ for a fixed pump-probe delay ($\Delta t_{\text{p1-pr}}=100\,\text{ps}$) with both pump pulses polarized vertically relative to the $ac$-plane of the sample (see Fig. S1).
To suppress coherent artifacts caused by interference of the two pump pulses near $\Delta t_{\text{p1-p2}}=0$, all actual measurements are performed with cross-polarized p1 and p2. Cross-correlation scans of the SHG signal as a function of the relative polarization angle $\alpha_{\text{p1,p2}}$ confirm the absence of measurable coherent artifacts at $\alpha_{\text{p1,p2}}=90^{\circ}$ (see Fig. S1). During double-pulse measurements, the SHG signal is recorded as a function of $\Delta t_{\text{p1-p2}}$ at each time point along the pump–probe delay, resulting in the two-dimensional data sets shown in Figs. 2c and 3a.
Analysis of the pump-probe data. For the time-resolved polarizer scan in Fig. 1d, the SHG intensity $I_{\text{SHG}}(\Delta t_{\text{p-pr}}\,\phi)$ is normalized to the average pre-pump value $I_{\text{SHG}}(\Delta t_{\text{p-pr}}<0,\phi=90\,^{\circ})$. For the fluence- and delay-dependent traces in Fig. 1e, the SHG intensity $I_{\text{SHG}}(\Delta t_{\text{p-pr}})$ is normalized to the average pre-pump value $I_{\text{SHG}}(\Delta t_{\text{p-pr}}<0)$. For clarity, traces corresponding to different fluences are vertically offset by a constant. To extract the fluence dependent amplitudes, decay constants, sliding mode frequencies and long-lived signal suppression (see Fig. 1f and 1g), the model function
$I_{\text{SHG}}(\Delta t_{\text{p-pr}})$ $=\left[\Delta t_{\text{p-pr}}\leq 0\right]+\left(1-\Delta I_{\text{PT}}-\Delta I_{\text{SP}}-\Delta I_{\text{fast}}^{\text{exc}}\cdot\Delta I_{\text{fast}}^{\text{decay}}\right)$ (2)
with the individual terms
$\Delta I_{\text{PT}}(\Delta t_{\text{p-pr}})$ $=\Delta I_{\text{PT}}^{0}\cdot\left(1-\text{exp}\left(-\frac{\Delta t_{\text{p-pr}}}{\tau_{\text{PT}}}\right)\right),$
$\Delta I_{\text{SP}}(\Delta t_{\text{p-pr}})$ $=-\Delta I_{\text{SP}}^{0}\cdot\text{cos}\left(\frac{2\pi\Delta t_{\text{p-pr}}}{T_{\text{sliding}}}+\phi_{\text{sliding}}\right)\cdot\text{exp}\left(-\frac{-\Delta t_{\text{p-pr}}}{\tau_{\text{sliding}}}\right),$
$\Delta I_{\text{fast}}^{\text{exc}}(\Delta t_{\text{p-pr}})$ $=\Delta I_{\text{fast}}^{\text{exc},0}\cdot\left(1-\text{exp}\left(-\frac{\Delta t_{\text{p-pr}}}{\tau_{\text{fast}}^{\text{exc}}}\right)\right),\text{ and}$
$\Delta I_{\text{fast}}^{\text{decay}}(\Delta t_{\text{p-pr}})$ $=\Delta I_{\text{fast}}^{\text{decay},0}\cdot\left(\text{exp}\left(-\frac{\Delta t_{\text{p-pr}}}{\tau_{\text{fast}}^{\text{decay}}}\right)\right)$
is fitted to the delay-dependent traces in Fig. 1d. Here, $\Delta t_{\text{p-pr}}$ is the pump-probe delay, $\Delta I_{\text{PT}}$ corresponds to the change in the SHG signal due to the shift of the equilibrium position after photo-excitation (see also Ref. [21] for details), $\Delta I_{\text{SP}}$ is the modulation of the SHG intensity due to coherent sliding motion, and $\Delta I_{\text{fast}}^{\text{exc}}$ or $\Delta I_{\text{fast}}^{\text{decay}}$, respectively, model the fast initial suppression of the SHG signal. Parameters $\tau_{\text{PT}}$ and $\tau_{\text{sliding}}$ correspond to the characteristic time scale of the structural shift or the sliding mode coherence time, respectively. Furthermore, $\tau_{\text{fast}}^{\text{exc}}$ and $\tau_{\text{fast}}^{\text{decay}}$ are the time constants for the excitation and subsequent decay of the fast initial signal suppression observed in the $\chi_{zyy}$ component. $T_{\text{sliding}}$ is the sliding mode period.
Analysis of the pump-pump-probe data. To determine the sliding mode amplitude $A_{\text{s}}$, frequency $f_{\text{s}}$, phase $\phi_{\text{s}}$, and damping $\gamma_{\text{s}}=\tau_{\text{s}}^{-1}$, as well as the quasi-dc signal suppression $\Delta I_{\text{const}}$ in double-pulse experiments, we fit the model
$I_{\text{SHG}}(\Delta t_{\text{p-pr}})=A_{\text{s}}\cdot\text{exp}\left(-\frac{\Delta t_{\text{p-pr}}}{\tau_{\text{s}}}\right)\cdot\text{cos}(2\pi f_{\text{s}}\Delta t_{\text{p-pr}}+\phi_{\text{s}})+\Delta I_{\text{const}}$ (3)
to the trSHG trace after the second excitation for each double-pulse delay $\Delta t_{\text{p-p}}$. The results are shown in Fig. 4a. The $\Delta t_{\text{p1-p2}}$-dependent sliding mode amplitude $A_{\text{s}}$ in Fig. 3f, on the other hand, is fitted by the model
$A_{\text{s}}(\Delta t_{\text{p1-p2}})=A_{\text{s}}^{\text{osc}}(\Delta t_{\text{p1-p2}})\cdot\left(1-A_{\text{e}}^{\text{damp}}\cdot\text{exp}\left(-\frac{\Delta t_{\text{p1-p2}}}{\tau_{\text{e}}^{\text{damp}}}\right)\right),$ (4)
where
$A_{\text{s}}^{\text{osc}}(\Delta t_{\text{p1-p2}})=A_{\text{s}}\sqrt{1+2\zeta\cdot(2\pi f_{\text{s}}\Delta t_{\text{p1-p2}})+\zeta^{2}}.$ (5)
Here, $A_{\text{e}}^{\text{damp}}$ is the leverage of electronic excitation for the damping of the sliding mode amplitude and $\tau_{\text{e}}^{\text{damp}}$ the electronic relaxation time. To extract the Fourier amplitudes at frequencies $f_{\text{s}}$ and $2f_{\text{s}}$, we multiply the delay-dependent SHG trace after the second excitation pulse with a super-Gaussian window function
$F_{\text{filt},t}=\text{exp}\left(-\left(\frac{(\Delta t_{\text{p1-p2}}-t^{\prime})^{2}}{2\sigma_{t}}\right)^{3}\right),$ (6)
before applying a standard fast Fourier transform algorithm. The quotient $(\Delta\gamma_{\text{s}}/\Delta A_{\text{s}})(T)$ is determined by fitting the linear model
$\gamma_{\text{s}}(A_{\text{s}})=A_{\text{s}}^{0}+(\Delta\gamma_{\text{s}}/\Delta A_{\text{s}})\cdot A_{\text{s}}$ (7)
to the two datasets presented in Fig. 4b.
Fitting of the polarization-dependent trSHG data. To fit the SHG polarimetry data shown in Fig. 1d, we consider the symmetry of the Td polytype. In the orthorhombic Td phase, WTe2 belongs to the $mm2$ point group. We consider only electric-dipole–allowed second-harmonic generation (SHG), described by the odd-parity, polar (i-type) second-order nonlinear susceptibility tensor $\chi^{(2)}$. According to Ref. [58], the non-vanishing components of $\chi^{(2)}$ for this symmetry are $\chi^{(2)}_{zzz}$, $\chi^{(2)}_{xxz}$, $\chi^{(2)}_{xzx}$, $\chi^{(2)}_{zxx}$, $\chi^{(2)}_{yyz}$, $\chi^{(2)}_{yzy}$, and $\chi^{(2)}_{zyy}$. The incident electric field is given by
$\bm{E}=E_{0}\begin{pmatrix}\cos(\varphi)/\sqrt{2}\\ \sin(\varphi)\\ \cos(\varphi)/\sqrt{2}\end{pmatrix}$, (8)
where $\varphi=0^{\circ}$ corresponds to p-polarized and $\varphi=90^{\circ}$ to s-polarized light relative to the plane of incidence. While SHG is collected in reflection at a $45^{\circ}$ angle with respect to the surface normal, for simplicity we perform the symmetry analysis in a transmission geometry, which yields equivalent selection rules. The induced second order nonlinear polarization $\bm{P}(2\omega)$ is thus given by
$\bm{P}(2\omega)$ $=\begin{pmatrix}2\chi_{xxz}E_{x}E_{z}\\ 2\chi_{yyz}E_{y}E_{z}\\ \chi_{zxx}E_{x}^{2}+\chi_{zyy}E_{y}^{2}+\chi_{zzz}E_{z}^{2}\end{pmatrix}$ (9)
$=E_{0}^{2}\begin{pmatrix}\chi_{xxz}\cos^{2}(\varphi)\\ \sqrt{2}\chi_{yyz}\sin(\varphi)\cos(\varphi)\\ \frac{1}{2}(\chi_{zzz}+\chi_{zxx})\cos^{2}(\varphi)+\chi_{zyy}\sin^{2}(\varphi)\end{pmatrix}$. (10)
Using trigonometric identities we yield
$\bm{P}(2\omega)=E_{0}^{2}\begin{pmatrix}\chi_{xxz}(1+\cos(2\varphi))/2\\ \chi_{yyz}\sin(2\varphi)/\sqrt{2}\\ \frac{1}{4}(\chi_{zzz}+\chi_{zxx})(1+\cos(2\varphi))+\frac{1}{2}\chi_{zyy}(1-\cos(2\varphi))\end{pmatrix}$. (11)
To calculate the SHG signal after the analyzer, we project the nonlinear polarization $\mathbf{P}(2\omega)$ onto the s- and p-polarized components using the unit vectors $\hat{\mathbf{v}}=(0,1,0)$ and $\hat{\mathbf{h}}=\frac{1}{\sqrt{2}}(1,0,1)$, respectively:
$I_{h}\sim|\hat{\bm{h}}\cdot\bm{P}|^{2}$ $=|\frac{1}{\sqrt{2}}(P_{x}+P_{z})|^{2}$ (12)
$=\frac{E_{0}^{4}}{2}|\chi_{\text{eff}}\cos^{2}(\varphi)+\chi_{zyy}\sin^{2}(\varphi)|^{2};$ (13)
$I_{v}\sim|P_{y}|^{2}$ $=2E_{0}^{4}|\chi_{yyz}\sin(\varphi)\cos(\varphi)|^{2},$ (14)
where $\chi_{\text{eff}}=\left(\chi_{xxz}+1/2(\chi_{zzz}+\chi_{zxx})\right)$. For $I_{h}$, we additionally assume a phase difference $\theta$ between $\chi_{\text{eff}}$ and $\chi_{zyy}$, leading to
$I_{h}\sim|\chi_{\text{eff}}\cos^{2}(\varphi)+e^{i\theta}\chi_{zyy}\sin^{2}(\varphi)|^{2}.$ (15)
Estimation of the transient lattice temperature increase. The transient temperature increase $\Delta T$ within the probed sample volume induced by the pump pulses in the pump–pump–probe experiments is estimated as
$\Delta T\approx\frac{F(1-R)}{\rho\,c_{\text{p}}\,d},$ (16)
where $F$ is the incident laser fluence, $R=0.49$ the reflectivity coefficient of WTe2 at the pump wavelength and normal incidence, $\rho=9.43\cdot 10^{3}\,\text{kg}\,\text{m}^{-3}$ is the material density, $c_{\text{p}}=1.77\cdot 10^{2}\,\text{J}\,\text{kg}^{-1}\,\text{K}^{-1}$ is the specific heat capacity at $300\,\text{K}$, $d=-\ln(1/\text{e})/\alpha=48\,\text{nm}$ is the approximate optical penetration depth with $\alpha=2.07\cdot 10^{5}\,\text{cm}^{-1}$ denoting the absorption coefficient at the pump wavelength. The penetration depth of the probe pulses ($\lambda_{\text{pr}}=800\,\text{nm}$, $d_{\text{pr}}=35\,\text{nm}$) is smaller than that of the pump pulse ($\lambda_{\text{p}}=1500\,\text{nm}$, $d_{\text{p}}=48\,\text{nm}$). We therefore neglect any effects arising from inhomogeneous excitation within the probed sample volume.
Possible origin of the $2\omega_{\text{s}}$ component. The appearance of sidebands in the Fourier transform of the time-dependent SHG signal can be understood within the framework of the deformation potential formalism. Following the arguments of Fukuda et al. [37], who observed similar features in linear optical pump–probe experiments on Td-MoTe2, we consider an optical transition near resonance between valence and conduction band states. In the presence of a coherent phonon, the corresponding optical susceptibility is given by
$\chi_{0}(\omega)=\frac{|P_{\text{cv}}|^{2}}{\varepsilon_{0}}\frac{1}{\hbar(\omega-(\omega_{\text{cv}}-\Delta\omega_{\text{DP}})+\text{i}\Gamma}.$ (17)
Here, $\hbar\omega$ is the photon energy, $\hbar\omega_{\text{cv}}$ the energy difference between valence and conduction band states, $\hbar\Delta\omega_{\text{DP}}$ the phonon-induced energy shift, $\varepsilon_{0}$ the vacuum permittivity, $P_{\text{cv}}$ the optical transition matrix element, and $\Gamma$ accounts for damping. The deformation potential term can be expressed as $\hbar\Delta\omega_{\text{DP}}=h_{0}Q(t)$, where $h_{0}$ reflects the electron–phonon coupling strength and $Q(t)$ is the time-dependent phonon amplitude. Assuming $h_{0}Q(t)\ll\hbar(\omega-\omega_{\text{cv}})+\mathrm{i}\Gamma$, we expand $\chi(\omega)$ in powers of $Q(t)$:
$\chi(\omega,Q(t))$ $=\chi_{0}\left(1+\frac{h_{0}Q(t)}{\hbar(\omega-\omega_{\text{cv}})+\text{i}\Gamma}\right)^{-1}$ (18)
$\approx\chi_{0}\left(1+\left(\frac{h_{0}}{\hbar(\omega-\omega_{\text{cv}})+\text{i}\Gamma}\right)Q(t)+\left(\frac{h_{0}}{\hbar(\omega-\omega_{\text{cv}})+\text{i}\Gamma}\right)^{2}Q(t)^{2}\right).$ (19)
For $Q(t)\sim Q_{0}\cos(\omega_{\text{s}}t)$, $Q(t)^{2}\sim Q_{0}^{2}\cos^{2}(\omega_{\text{s}}t)=(Q_{0}^{2}/2)(1+\cos(2\omega_{\text{s}}t))$. With this, the second-order nonlinear polarization is given by
$P^{2\omega}_{i}$ $=\varepsilon_{0}\chi_{ijk}(\omega,Q(t))E_{j}E_{k}$ (20)
$=\varepsilon_{0}\chi_{ijk}^{0}(\omega)\left(1+\frac{Q_{0}^{2}\beta^{2}}{2}+Q_{0}\beta\cos(\omega_{\text{s}}t)+Q_{0}^{2}\beta^{2}\cos(2\omega_{\text{s}}t)\right)E_{j}E_{k},$ (21)
with $\beta=h_{0}/\hbar(\omega-\omega_{\text{cv}})+\text{i}\Gamma$, resulting in oscillations at frequency $2\omega_{\text{s}}$ of the SHG intensity in the time-domain.
Computational details. Density functional theory calculations are conducted within the plane-wave pseudopotential code Quantum ESPRESSO [59, 60], using fully-relativistic norm-conserving Pseudopotentials [61] and the Perdew-Burke-Ernzerhof generalized-gradient approximation (PBE) for the exchange-correlation functional [62]. The plane-wave kinetic-energy cutoff is set to 80 Ry, the Brillouin zone is sampled with a $10\times 6\times 4$ Monkhorst-Pack grid and the phonon dispersion is obtained from density functional perturbation theory on a $5\times 3\times 2$ q-point mesh [63]. The crystal lattice vectors are chosen to reproduce the experimental one from Ref. [64], and the crystal structure is relaxed within that unit cell. The electron and phonon eigenvalues, as well as the electron-phonon coupling matrix $g_{nm}^{\rm\nu}({\bf k},{\bf q})$, are interpolated on a dense $40\times 24\times 16$ grid via maximally-localized Wannier functions [65] within the EPW code [66], which uses Wannier90 as a library [67].
Ab initio description of coherent phonon excitation. The theoretical framework for the simulation of the coherent lattice dynamics is based on the solution of the coherent phonon equation of motion [38]:
$\partial_{t}^{2}Q_{\nu}+\gamma_{\nu}\partial_{t}Q_{\nu}+\omega^{2}_{\nu}Q_{\nu}=D_{\nu}(t)\quad.$ (22)
Here, $\partial_{t}=\partial/\partial_{t}$ and $Q_{\nu}$ is the adimensional displacement amplitude of the coherent phonon with frequency $\omega_{\nu}$. $\gamma_{\nu}$ is the coherent phonon damping rate mediated by phonon-phonon interactions [39]. The room-temperature coherent phonon damping rate is estimated from the third-order force constant, which is obtained from finite differences using the third-order.py utility of the shengBTE code [68]. The term on the right-hand side of Eq. 22 is the coherent phonon driving force due to electron-phonon interaction:
$D_{\nu}(t)=-\frac{2\omega_{\nu}}{\hbar}\sum_{n\mathbf{k}}g_{nn}^{\nu}(\mathbf{k},0)\Delta f_{n\mathbf{k}}(t)\quad,$ (23)
where $\Delta f_{n{\bf k}}=f_{n{\bf k}}(t)-f_{n{\bf k}}^{(0)}$ is the change of the occupation function with $f_{n{\bf k}}^{(0)}$ the (equilibrium) Fermi-Dirac distribution before excitation. The time-dependent electron distribution function is obtained within the time-dependent Boltzmann equations (TDBE) [69, 70] as implemented in the EPW code. The TDBE is solved using an excited-state Ansatz for the electrons described by a high-temperature Fermi-Dirac distribution $f_{n{\bf k}}(t=0)=[{\rm exp}((\varepsilon_{n{\bf k}}-\mu)/k_{\rm B}T_{\rm exc})+1]^{-1}$, where the excitation temperature $T_{\rm exc}$ is set to match the energy transferred to the system by an incident pump with fluence $F$, as discussed in Ref. [38]. The TDBE are solved through second-order Runge-Kutta time-stepping with a time interval of 1 fs, within an energy window of 4 eV around the Fermi energy. The coherent phonon amplitude obtained from Eqs. (22) and (23) is illustrated in Fig. S3 (a) (single-pulse excitation).
Based on this framework, the maximum coherent-phonon amplitude $Q_{\nu}^{\rm max}(F)$ is estimated by considering the static limit of Eq. 22 ($\partial_{t}Q^{\nu}=0$) leading to $Q_{\nu}^{\rm max}(F)=D_{\nu}/\omega^{2}_{\nu}$. This value constitutes an upper theoretical limit to the maximum lattice displacement that the crystal can undergo for a given electronic excitation $\Delta f_{n\mathbf{k}}$.
To simulate the coherent lattice dynamics induced by two distinct laser pulses, the coherent phonon equation of motion is solved by considering the action of two driving forces:
$D^{\nu}_{\rm dp}(t;\Delta t_{p_{1}-p_{2}})=D_{\nu}(t)+D_{\nu}(t-\Delta t_{p_{1}-p_{2}})\theta(\Delta t_{p_{1}-p_{2}})\quad,$ (24)
where $\Delta t_{p_{1}-p_{2}}$ is the time delay between the two pump pulses. The coherent phonon amplitudes obtained from the time propagation of Eqs. ((22)-(24)) is illustrated in Figs. S3 (b-d) for different values of the time delay $\Delta t_{p_{1}-p_{2}}$ between the two pulses. The frequency renormalization proportional to the coherent phonon amplitude observed in experiments can be phenomenologically accounted for by substituting the bare phonon frequency $\omega_{\nu}$ with the dressed frequency $\Omega_{\nu}=\omega_{\nu}-\beta|Q_{\nu}|$ in Eq. (22), where $\beta$ is a positive fitting parameter.
Declarations
  • Acknowledgements. This work was funded by an ETH Postdoctoral Fellowship and a Swiss National Science Foundation (SNSF) Postdoctoral Fellowship (TMPFP2 217303). The authors would like to thank M. Bauer, A. Lindenberg, S. L. Johnson, D. Soranzio, M. Savoini, and K. S. Burch for helpful discussions as well as J. Hecht and S. Reitz for technical support.
  • Author contributions. The project was conceived by J.G.H. Experiments and data analysis were conducted by J.G.H., with contributions from A.C. and T.L. Q.M. carried out the symmetry analysis and related DFT calculations. C.E. and F.C. performed the DFT calculations on the energy-dependent electron-phonon coupling. The manuscript was written by J.G.H. M.F. provided overall guidance and support. All authors discussed the results and commented on the manuscript.
  • Competing interests. The authors declare no competing interests.
  • Materials & Correspondence. Correspondence and requests for materials should be addressed to J.G.H.
  • Data availability. The data that support the findings of this study are available from the corresponding author upon reasonable request.
  • Code availability. The code supporting the findings of this study is available from the corresponding author upon reasonable request.
Figure 1: $\mid$ Ultrafast structural dynamics in ferroelectric Td-WTe2. a, Schematic of the experimental setup. b, Structural configurations of WTe2 and lattice modes connecting the Td and 1TT’ polymorphs to the paraelectric high-symmetry phase. Grey background structures denote the atomic configuration of the high-symmetry paraelectric phase, included for comparison. c, Illustration of structural dynamics on the excited-state potential energy surface following photo-excitation. d, Polarization-dependent SHG intensity at selected pump–probe delays $\Delta t_{\text{p-pr}}$. The incident polarization of the fundamental is varied, with the analyzer fixed in the horizontal orientation. e, Normalized, delay-dependent SHG efficiency (polarizer: vertical; analyzer: horizontal) as a function of absorbed laser fluence $F$. Traces are vertically offset for clarity. f, Sliding mode amplitude $A_{\text{s}}$ and lifetime $\tau_{\text{s}}$ as a function of absorbed fluence $F$. g, Sliding mode frequency and quasi-static suppression of SHG intensity as functions of absorbed fluence $F$.
Figure 2: $\mid$ Ab initio modeling of sliding amplitude saturation. a, State-resolved electron-phonon coupling strength at the sliding mode frequency at two electronic temperatures $T_{\text{elec}}=2000\,\text{K}$ and $T_{\text{elec}}=4000\,\text{K}$. Red (blue) indicates positive (negative) contributions to sliding mode excitation with more saturated colors corresponding to stronger coupling. Along the energy axis, the coupling strength is weighted by the change in the Fermi–Dirac distribution $\Delta f_{\text{elec}}$ at the respective temperature. b, Calculated maximum amplitudes $Q_{\text{max}}$ of the sliding mode and two higher-frequency modes of Td-WTe2 as a function of absorbed fluence $F$. Triangular data points represent experimentally determined sliding mode amplitudes for comparison.
Figure 3: $\mid$ Coherent control and amplification of sliding motion. a, Schematic of amplitude control of displacively excited coherent phonons using double-pulse excitation. b, Normalized SHG intensity as a function of pump–probe delay $\Delta t_{\text{p-pr}}$ and double-pulse delay $\Delta t_{\text{p1-p2}}$ for a combined absorbed fluence of $2.1\,\text{mJ}\,\text{cm}^{-2}$. Red, yellow, and violet arrows indicate the double-pulse delays corresponding to the pump–probe traces shown in c. c, Pump–probe SHG traces at selected double-pulse delays, demonstrating coherent control of the sliding mode amplitude. d, Simulated sliding mode amplitude $Q$ as a function of $\Delta t_{\text{p-pr}}$ and $\Delta t_{\text{p1-p2}}$, assuming an amplitude-dependent sliding frequency.e, Comparison of pump–probe SHG traces for $\Delta t_{\text{p1-p2}}=0\,\text{ps}$ and $\Delta t_{\text{p1-p2}}=8.3\,\text{ps}$. f, Sliding mode amplitude after the second optical excitation as a function of $\Delta t_{\text{p1-p2}}$. Dashed black line, expected behavior in coherent control experiments without amplitude saturation (see Methods Section); red line, fitted model incorporating amplitude suppression at short double-pulse delays; violet line, exponential component of the model used to extract the timescale of transient sliding mode suppression. g, Comparison of vibrational amplitudes extracted from pump–probe ($P$) and pump–pump–probe (PPP) measurements. Black line, estimated vibrational amplitude assuming a linear dependence of $A_{\text{s}}$ on the excited carrier density.
Figure 4: $\mid$ High-amplitude vibrational spectroscopy in targeted excited states. a, Sliding mode amplitude $A_{\text{s}}$, damping $\gamma_{\text{s}}=\tau_{\text{s}}^{-1}$, and frequency $f_{\text{s}}$ as a function of the double-pulse delay $\Delta t_{\text{p1-p2}}$ for a combined absorbed fluence of $2.1\,\text{mJ}\,\text{cm}^{-2}$. Light-colored shades, $1\sigma$-confidence intervals of the fits used to extract parameters from the two-dimensional dataset shown in Fig. 3b. b, Sliding mode damping constant $\gamma_{\text{s}}$ as a function of the sliding mode amplitude $A_{\text{s}}$ for two different combined fluences in double-pulse experiments. Black line and markers, values obtained in pump-probe experiments (see Fig. 1f); solid red and violet lines, linear models fitted to the data to determine $(\Delta\gamma/\Delta A)_{\text{s}}$; errorbars, $1\sigma$-confidence intervals of the fits used to extract $A_{\text{s}}$ and $\gamma_{\text{s}}$. c, Low-energy phonon band structure of WTe2.
Figure 5: $\mid$ Modulation of nonlinear optical properties by high-amplitude coherent sliding motion. a, Fourier-transform of the two-dimensional dataset in Fig. 3b along the $\Delta t_{\text{p-pr}}$ axis. The data has been normalized to the maximum spectral density at the fundamental sliding mode frequency. b, Fourier spectra extracted from a at three selected double-pulse delays. c, The $2\omega_{\text{s}}$-contribution to the SHG signal as a function of the double-pulse delay $\Delta t_{\text{p1-p2}}$. The fundamental sliding mode amplitude determined from fits (see Figs. 3 and 4) is shown for comparison. d, Sliding-mode-amplitude-dependent $2\omega_{\text{s}}$-contribution.
Fig. S1 $\mid$ Polarization-dependent SHG signal in double-pulse and single-pulse experiments a, Comparison of SHG signals from WTe2 as a function of the delay between two pump pulses for parallel (violet) and cross-polarized (red) configurations (combined absorbed fluence $F=2.1\,\text{mJ}\,\text{cm}^{-2}$). The probe pulse interrogates the sample at a fixed delay of $\Delta t_{\text{p-pr}}=100\,\text{ps}$, where coherent phonon oscillations are negligible. For parallel-polarized pump pulses, interference at the sample surface leads to a delay-dependent modulation of energy absorption, resulting in a corresponding variation of the SHG signal. In contrast, interference effects are strongly suppressed for cross-polarized pump pulses. Consequently, all double-pulse excitation experiments were performed using cross-polarized pump pulses. b, SHG intensity measured at $\Delta t_{\text{p-pr}}=100\,\text{ps}$ for two different polarizer–analyzer configurations, shown as a function of the polarization angle of a single pump pulse (absorbed fluence $F=1.6\,\text{mJ}\,\text{cm}^{-2}$; $0^{\circ}:P\parallel a$-axis; $90^{\circ}:P\parallel b$-axis of the crystal). Only a weak anisotropy is observed, indicating that in double-pulse experiments both pump pulses, despite their differing polarizations, contribute similarly to the electronic excitation of the material.
Fig. S2 $\mid$ Time-dependent polarimetry measurements a–d, SHG intensity at selected pump–probe delays $\Delta t_{\text{p-pr}}$ as a function of the incident polarization angle of the fundamental light. The analyzer was fixed in the horizontal direction. Violet circles, experimental data; violet lines, fits based on the material’s symmetry (see Methods Section for details). e, Comparison of polarization-resolved SHG intensity just before and immediately after time zero. A pronounced change is observed in the $\chi_{zyy}$ component (at $90^{\circ}$), while the $\chi_{\text{mix}}$ component (at $0^{\circ}$) remains largely unaffected. f, Delay-dependent evolution of the SHG tensor components $\chi_{zyy}$ (violet) and $\chi_{\mathrm{mix}}$ (red), extracted from fits to the time-resolved polarimetry data. Upon photoexcitation, $\chi_{zyy}$ exhibits a rapid drop to a quasi-static level with superimposed oscillations at the sliding-mode frequency. In contrast, $\chi_{\mathrm{mix}}$ shows neither a fast initial suppression nor a comparable long-lived change.
Fig. S3 $\mid$ Time-resolved simulations of the coherent phonon excitation. a Time-dependent coherent phonon amplitude $Q_{\nu}$ for a single driving pulse, obtained from the solution of Eqs. ((22)-(23)), considering a single driving force marked in brown. b-c Coherent phonon amplitude and driving force in the presence of two-pulse excitation as described via Eqs. ((22)-(24)) for different time delays $\Delta t_{p_{1}-p_{2}}$ between the two pump pulses. Panel b illustrates the case in which the time delay coincides with one full phonon period, resulting in resonant excitation. In panels c and d, the time delay is set to 1.25 and 1.5 phonon periods. In these cases, the second pulse results in the destructive interference of the coherent phonon, as revealed by the lower coherent phonon amplitude compared to the resonant case.
Fig. S4 $\mid$ Double-pulse excitation of coherent phonons. Coherent phonon amplitude as a function of time delay between two pump pulses $\Delta t_{p_{1}-p_{2}}$ and of the pump-probe delay $\Delta t_{p_{2}-p_{r}}$. a Coherent phonon amplitude obtained from Eq. 22 in absence of damping ($\gamma_{\nu}=0$). b Coherent phonon amplitude obtained from Eq. 22 in presence of damping. c Coherent phonon amplitude obtained by considering the frequency renormalization according to $\Omega_{\nu}=\omega_{\nu}-\beta|Q_{\nu}|$ in Eq. (22).